# Multi-Lattice Sampling of Quantum Field Theories via Neural Operators

### 01 Jan 2024 v1

@@ 1 Introduction
Let $S$ an action on some lattice (as an approximation of the continuous action $\mathcal{S}$). Albergo et al. [1] propose to sample from the quantum field theory defined by $S$ by using a normalizing flow to parametrise a density $q_{\theta}$ on the lattice and optimize its parameters $\theta$ until $q_{\theta}$ is a good approximation of $Z^{-1}e^{-S}$. In particular, such flow models are inherently tied to the chosen lattice as there is no easy way of using them given a different choice of lattice.
On the other hand, operator learning promotes the viewpoint that the lattice/mesh is merely a computational tool, and model should capture the underlying continuous physics. Kovachki et al. [2] term this property of models discretization invariance.  [^1] Discretization invariance means that the neural operator evaluated on finer and finer discretizations approximates the continuous operator. Thus, strictly speaking, it is not a requirement of invariance rather that of convergence.
@@ 2.3 Neural Operators
Neural operators[2] are trainable function-to-function mappings. In particular, both their domains and codomains are infinite dimensional function spaces. In practice, one works with neural operators by choosing a mesh/lattice $X\subset\mathbb{R}^{n}$, representing functions by their evaluations on $X$ and let the neural operator operate on this collection of evaluations. By design, neural operators can be evaluated on lattices of different size. Importantly, if a neural operator is applied to a sequence of meshes $X_{i}$, approaching the continuum limit $X_{i}\rightarrow\mathbb{R}^{n}$, it converges to the underlying continuous operator. The main use case of neural operators is to approximate the solution of partial differential equations, i.e. learn the mapping from an initial condition to the time evolved state after some time $\Delta t$ (Figure 2). We will use them for parametrizing a flow, i.e. a vector field $\mathcal{V}_{t}$ connecting the free theory (base density) to the $\phi^{4}$-theory (target density) in a way that can be evaluated at any mesh.
@@ 3 Flows parametrised by neural operators
    Use a per-node neural network $f_{\theta}$ to embed the field values, $\phi_{emb}=f_{\theta}(\phi)\in\mathbb{R}^{c\times N\times....\times N}$
@@
    Apply a per-node neural network $\tau_{\theta}$ to the concatenation $(C,\phi_{emb})$ with output $Y=\tau_{\theta}(C,\phi_{emb})\in\mathbb{R}^{T\times N\times....\times N}$

### 13 Jan 2024 v2

# Multi-Lattice Sampling of Quantum Field Theories via Neural Operator-based Flows

@@ 1 Introduction
Let $S$ an action on some lattice (as an approximation of the continuous action $\mathcal{S}$). Albergo et al. [1] propose to sample from the quantum field theory defined by $S$ by using a normalizing flow to parametrise a density $q_{\theta}$ on the lattice and optimize its parameters $\theta$ until $q_{\theta}$ is a good approximation of $Z^{-1}e^{-S}$.
Operator learning promotes the viewpoint that the lattice/mesh is merely a computational tool, and model should capture the underlying continuous physics. Kovachki et al. [2] term this property of models discretization invariance.  [^1] Discretization invariance means that the neural operator evaluated on finer and finer discretizations approximates the continuous operator. Thus, strictly speaking, it is not a requirement of invariance rather that of convergence.
@@ 2.3 Neural Operators
Neural operators[2] are trainable function-to-function mappings. In particular, both their domains and codomains are infinite dimensional function spaces. In practice, one works with neural operators by choosing a mesh/lattice $X\subset\mathbb{R}^{n}$, representing functions by their evaluations on $X$ and let the neural operator operate on this collection of evaluations. By design, neural operators can be evaluated on lattices of different size. Importantly, if a neural operator is applied to a sequence of meshes $X_{i}$, approaching the continuum limit $X_{i}\rightarrow\mathbb{R}^{n}$, it converges to the underlying continuous operator. The main use case of neural operators is to approximate the solution of partial differential equations, i.e. learn the mapping from an initial condition to the time evolved state after some time $\Delta t$ (Figure 2), but have also been applied for multi-resolution generative modelling [18, 19]. We will use them for parametrizing a flow, i.e. a vector field $\mathcal{V}_{t}$ connecting the free theory (base density) to the $\phi^{4}$-theory (target density) in a way that can be evaluated at any mesh.
@@ 3 Flows parametrised by neural operators
    Use a per-node neural network $f_{\theta}$ to embed the field values, $\phi_{emb}=f_{\theta}(\phi)\in\mathbb{R}^{c\times N\times....\times N}$.
@@
    Apply a per-node neural network $\tau_{\theta}$ to the concatenation $(C,\phi_{emb})$ with output $Y=\tau_{\theta}(C,\phi_{emb})\in\mathbb{R}^{T\times N\times....\times N}$.